Predict week 11 using the different methods requested and then use historical analysis to determine the best method.

Smoothing:
Week Attendance

1 350
2 500
3 750
4 705
5 820
6 743
7 880
8 900
9 795
10 900 three week four week three week weighted

Next Week Projection (week 11):

Longest Regression:

1) Use regression analysis to predict week 11 in the above example:

Regression can also be used for predictions within the relevant range:
1
Expert's answer
2019-09-23T07:05:11-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Week & Attendance \\ \hline
 1 & 350 \\
 \hdashline
2 & 500 \\
 \hdashline
3 & 750 \\
 \hdashline
4 & 705 \\
 \hdashline
5 & 820 \\
 \hdashline
6 & 743 \\
 \hdashline
7 & 880 \\
 \hdashline
8 & 900 \\
 \hdashline
 9 & 795\\
 \hdashline
10 & 900 \\
 \hline
\end{array}"

"\sum x_i=55, \sum y_i=7343"


"\sum x_i^2=385, \sum y_i^2=5682899"

"\sum x_iy_i=44493"



"mean:\ \bar{x}={\sum x_i \over n},\ \bar{y}={\sum y_i \over n}""mean:\ \bar{x}={55\over 10}=5.5,\ \bar{y}={7343 \over 10}=734.3""Trend \ line: y=A+Bx, B={S_{xy} \over S_{xx}}, A=\bar{y}-B\bar{x}""correlation\ coefficient: r={S_{xy}\over \sqrt{S_{xx}}\sqrt{S_{yy}}}"

"S_{xx}=\sum(x_i-\bar{x})^2=\sum x_i^2-n\cdot\bar{x}^2"

"S_{xx}=385-10\cdot5.5^2=82.5"

"S_{yy}=\sum(y_i-\bar{y})^2=\sum y_i^2-n\cdot\bar{y}^2"

"S_{yy}=5682899-10\cdot734.3^2=290934.1"


"S_{xy}=\sum(x_i-\bar{x})(y_i-\bar{y})=\sum x_iy_i-n\cdot\bar{x}\bar{y}"

"S_{xy}=44493-10\cdot5.5\cdot734.3=4106.5"


"B={S_{xy} \over S_{xx}}={4106.5\over 82.5}\approx49.77575758"

"A=\bar{y}-B\bar{x}=734.3-49.77575758\cdot5.5=460.5333333"


"r={S_{xy}\over \sqrt{S_{xx}}\sqrt{S_{yy}}}={4106.5\over \sqrt{82.5}\sqrt{290934.1}}\approx0.8382"

"y=460.5333333+49.77575758x"

"y(11)=460.5333333+49.77575758(11)\approx1008"